OpenStudy (acespeedfighter):

If 9x = 63, then x =_______ . (Only input whole number) (19 point)

OpenStudy (sql_server):

Consider this example. \[A*B=C\] \[B=\frac{ C }{ A }\] Can you try it now?

OpenStudy (acespeedfighter):

assuming A=9 and B=63 567

OpenStudy (sql_server):

check it again. You didn't do it right.

OpenStudy (acespeedfighter):

im confused

OpenStudy (sql_server):

see the sequence of the terms in the question and the example then try it.

OpenStudy (acespeedfighter):

hmmm

OpenStudy (acespeedfighter):

i dont know

OpenStudy (sql_server):

In the question, \[9*x=63\] In the example, \[A*B=C\] If \[B=\frac{ C }{ A }\] What would be the value of x?

OpenStudy (acespeedfighter):

7

OpenStudy (sql_server):

yes correct. Now remember this process throughout your lifetime so that you'll never get confused again.

OpenStudy (acespeedfighter):

yea got it :D